# Serializing Java Objects in Plain Code

### v3

@@ Abstract
[…]
mpdframe frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, innermargin =0.1cm, outermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm,
@@ 1 Introduction
However, readability is an aspect of serialization that has been neglected in previous work. Producing readable serialized objects is the problem we address in this paper. This is of utmost importance when serialized objects are consumed by humans. For example, when serialized objects are included in test cases as test inputs[10, 11, 12, 13], readability is essential to let developers understand the tests. Clearly, mainstream serialization techniques fall short with respect to readability. For example, 1 shows an example of a test case generated by the automated test generation tool Pankti [12], a tool that creates tests from objects at runtime. In 1, the test inputs are in the XML format, as is the case for the data used to populate the receiver and key variables. XML strings arguably makes the code hard to read, making it challenging to understand the test’s intention.
@@ 24 COSName key = deserializeFromXML(
To solve the problem of unreadable serialized data, we propose the novel concept of plain-code serialization, which consists in serializing an object directly into code statements. By serializing objects to plain code, the serialized representation becomes readable, using a syntax that is both concise and natural to developers. We illustrate this with the semantically equivalent test of 2, which contains the same serialized objects as 1 represented as plain code, instead of XML. In addition, the compiler gives us strong guarantees about the static correctness of the serialized objects. Deserialization then simply means executing the code. In the context of tests, this completely removes dirty data like XML strings, thus making the generated tests more readable and self-contained [17].
@@ 2.1 Concept & Overview
$reconstruct(x)=evaluate(plain\_code\_serialize(x))$ (2)
The equivalence relationship ($\equiv$) means that the object after reconstruction is in the same state as the one observed at runtime. While Equation 1 applies to any serialization library, Equation 2 is specific to plain-code serialization, with $reconstruct$ being the evaluation of the plain-code serialized representation of $x$ obtained from the function $plain\_\allowbreak{}code\_\allowbreak{}serialize$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 2.2 Structure-based Serialization (sb)
The core idea of the structure-based object serialization strategy is to statically analyze a class and build a reconstruction plan. This is done by finding actions which set the value of fields, such as constructor or setter method calls. The plan is instantiated at runtime for a given object.
@@ 2.2.2 Optimizing Readability
As is the case for Monkey, we enumerate all possible reconstruction plans for a target class. However, we are interested in the shortest and most idiomatic plan, to reconstruct an object using the minimum number of statements [18]. To quantify how idiomatic a reconstruction plan is, we first assign each action a pre-defined integer cost value. For example, calling a constructor has a very low cost, while directly setting a field has a higher cost. We then select the reconstruction plan with the lowest cumulative cost. We formulate this optimization as a constraint problem, shown in subsubsection 2.2.2. In the constructed problem, every action $i$ is assigned a boolean variable $a_{i}$. If an action sets a field $f$, it is added to the set $\beta_{f}$, and if it is a constructing action, also to $\alpha$.
@@ 2.3.1 Capturing Events
An event of interest is the occurrence of an operation which mutates an object. Events are constructor calls (ConstructEvent), field assignments (FieldSetEvent), or impure methods calls. Method call events are split into a start (MethodStartEvent) and an end event (MethodEndEvent), allowing other events to happen in-between. We capture all calls to methods with at least one field assignment.
We denote all potential events to reconstruct an object $o$ as $event\_sequence(o)$. Based on this notation, we specialize the general object reconstruction formulation, Equation 1, for trace-based reconstruction as follows:
@@
Given a type we want to serialize, we first identify all types needed for its reconstruction because they are used either as fields, constructor parameters, or method parameters. Then, we instrument the constructors, accessible methods, and field assignments for all these types. For example, in the Habitat class (3), we instrument the constructor on line $28$, as well as the grow method on lines $30-33$. We show the resulting instrumented code for the grow method in 6. When the instrumented grow is invoked at runtime, it emits multiple events, ordered by a monotonically increasing counter representing logical time.
@@ 2.3.3 Object Reconstruction Mixing
We show such a case in 8, where we are constructing a Monkey. Monkey is serialized using structure-based serialization, and its habitat field is serialized using trace-based serialization. As the captured events are mapped to code offline, but the plain-code serialized habitat field has to be inserted into a specific spot in the generated code, the structure-based serialization strategy uses the “use object reference” (from Table 1) marker action. This action emits an $id@logical-time$ reference, which is replaced by the trace-based serialized value after the event sequence conversion is finished. Besides this action, Table 1 also contains two more actions, “use static field” and “use standard charset”. These are actions that can not be done structurally, but also do not need the full trace-based serialization machinery. For the “use static field” action type, we look at all static fields of the class we want to serialize and detect if any of these store a reference-equal object. If such a field can be found, we emit a field read instead of tracing the constructor and subsequent mutating method calls. A similar, but more specific, optimization is realized by the “use standard charset” action type. This action type replaces references to Java’s standard charsets with the canonical named constants, and serves as an example of integrating domain logic into the serialization process. To further facilitate extensibility, such custom adapters and action types can be provided as plugins.
@@ 2.4.4 Implementation
ProDJ relies on Spoon [19] for source code analysis, in order to identify the associated types of methods involving serialization within the target Java application. ProDJ uses Z3 [20] to solve the constraint problem of subsubsection 2.2.2 and determine the most idiomatic reconstruction plan for types that are structure-based serializable. The event sequences for trace-based serializable objects are logged in JSON. The instrumentation of methods and their associated types is achieved using ByteBuddy [^3] https://bytebuddy.net/.
@@ 3.2 Subjects
Table 2: Projects used in the evaluation of plain code serialization: the exact version (SHA) we use for our experiments, the number of lines of Java code (LOC), the number of serialization points we consider (#SPs), the number of types related to the serialization points (#Types), the number of serialization triggered in production by our workloads (#SPs Prod), and the number of types that we target for plain code serialization at runtime (#Types Prod).
We evaluate ProDJ with $4$ open-source Java projects that are popular and active. We present them in Table 2. PDFBox is a library and command-line tool for the manipulation of PDF documents [^5] https://pdfbox.apache.org/. Its features include conversion between text and PDF documents, extraction of images, and merging of PDF documents, among others. GraphHopper is an open-source routing application [^6] https://www.graphhopper.com/, used to compute a path between input locations on a map. BroadLeaf is an e-commerce web application [^7] https://www.broadleafcommerce.com/. Gephi is a graphical tool for working with graph data [^8] https://gephi.org/, including capabilities for visualization and analysis.
For each project in Table 2, column $2$ links to the exact version (SHA) we use for our evaluation, and column $3$ presents the number of lines of Java source code within each project, calculated with tokei. Columns $4$ and $5$ give the number of methods instrumented by ProDJ within the project (#SPs for ‘serialization point’), and consequently, the number of associated types (#Types) that need to be instrumented at this point. For example, ProDJ adds serialization points in $2,933$ methods within BroadLeaf, involving $2,223$ types. For our evaluation, we select methods within each project that are public, non-abstract, non-static, non-deprecated, contain more than one statement, and are declared in a public, non-anonymous, non-local, and non-deprecated class. Columns $6$ and $7$ of Table 2 provide the same metrics, but computed after runtime observations, i.e., the number of serialization points and types that are covered at runtime, which we discuss in the next subsection.
@@ 3.3 Workloads
For our experiments, we design workloads that are typical for each of the $4$ projects. We invoke PDFBox through its command line interface, and use its “render” command to convert a PDF document containing multiple pages into JPEG image files. We run the GraphHopper server, and query GraphHopper for a route between two coordinates in Berlin [^9] https://bit.ly/3LA2PVu. In BroadLeaf, we add some products to our shopping cart, create a user account, log in, and favorite some products. Next, we browse through our shopping cart and change the quantity of some products, before finally completing our order. We use the graphical interface of Gephi to load the “Les Misérables” dataset [^10] https://github.com/gephi/gephi/wiki/Datasets#social-networks, and compute all the statistics on it that Gephi offers. This includes computing the average degree of nodes, the network diameter, as well as operations like connected components or Pagerank.
Our interactions with each project result in the invocation of serialization points and correspondingly, the instantiation of types that are instrumented. The last two columns of Table 2 give those numbers. For example, of the $2,933$ methods that are instrumented by ProDJ for BroadLeaf, $508$ are triggered by our workload (column #SPs Prod). Consequently, $444$ unique types involved in serialization are instantiated (column #Types Prod). As we shall see in subsection 4.1, these serialization points will trigger the serialization of thousands of objects under our workload.
@@ 4 Experimental Results
Table 3: Results of plain-code serialization on $4$ open-source, real-world Java applications
@@ 4.1 RQ1 (Objects)
{mdframed}
[nobreak=true,style=mpdframe,frametitle=Answer to RQ1] ProDJ successfully plain-code serializes thousands of objects ($174,699$) observed at runtime in $4$ real-world Java applications. Both the structure-based and trace-based serialization strategies are useful to translate runtime objects in plain Java.
@@
3// GHPoint uses the structure-based strategy
[…]
6// The point list uses the structure-based strategy
[…]
8object.setPoints(points1);
9$\cdots$
@@ 4.2 RQ2 (Accuracy)
[…]
@@
[…]
@@
[…]
@@
In 10, we present one test generated by ProDJ for GraphHopper, clearly following the Arrange-Act-Assert pattern. This test verifies the behavior of the alignOrientation method of the AngleCalc class in GraphHopper. ProDJ generates the Java statements from lines $5$ to $7$ at runtime using the structure-based strategy. Note that ProDJ reuses the parameter names of alignOrientation for the local variables in the test on lines $6$ and $7$. The assertion on line $11$ verifies that the double value output from the invocation of alignOrientation on line $9$ matches the value recorded at runtime. We also present one ProDJ-generated test for PDFBox in 11. The method under test, called getComponents defined in the class PDColor, returns a non-null, float array output, based on the value of a field called colorSpace. Note that ProDJ outlines the recreation of the receiver object of type PDColor into a separate createPDColor method. Within this method, ProDJ uses object reconstruction mixing for plain-code serializing PDColor. Specifically, during the pre-execution phase, ProDJ discovers that the associated type COSArray is eligible for trace-based serialization. At runtime, ProDJ records a constructor call event for COSArray (recreated on line $15$), followed by a call to the method addAll (line $22$). The only parameter of addAll is a list, which ProDJ serializes using the structure-based strategy (line $20$). However, the contents of this list are serialized using the trace-based strategy (lines $16$ $18$). ProDJ uses the parameter name objectList of addAll to derive the names of the local variables in the helper method. Note that ProDJ observes the calls to the factory method COSFloat.get at runtime and therefore recreates these calls, instead of invoking the COSBase constructor directly. The String arguments for these COSFloat.get method calls are directly inlined, in contrast with the double arguments on lines $6$ and $7$ in the first test of 10. This is because ProDJ tries to inline primitive values, but only if this would not break up the Arrange-Act-Assert pattern. Next, on line $23$ of 11, a PDColorSpace object is reconstructed using the runtime trace as being an instance of the RGB color space. Finally, the two-argument constructor call on line $24$ reuses the reconstructed objects to recreate the runtime PDColor object. This object is used within the generated test as the receiver for the invocation of getComponents (line $8$). Note also the inline creation of the float array on line $10$ within the assertion, which verifies the equality of the expected and actual float array output. We argue that these plain-code serialization and readability optimization strategies employed by ProDJ result in tests that not only reflect production observations, but are also focused, well-structured, an easy-to-read.
@@
All test cases generated by ProDJ compile correctly and when we execute them, $1,038$ test cases pass, which is a success rate of $59.9$%. Specifically, for three of the four case studies, a large majority of the generated tests pass. For PDFBox, $82$% of all tests generated pass, the highest among the four projects. That is a strong property. Given the structure of generated tests, it means that the state of an object after calling a method, i.e., after the ‘Act’ phase of the test, is consistent with the state observed in production (the ‘Assert’ phase). While not being a correctness proof, this is strong consistency evidence. In addition, manual analysis of a random sample of plain-serialized objects in tests, such as the one in 9, confirm that they are meaningful.
@@
{mdframed}
[nobreak=true,style=mpdframe,frametitle=Answer to RQ2] Our experiment with using plain-code serialization for testing shows that ProDJ faithfully recreates runtime states for all objects in nearly 60% of the generated tests. The assertions in this testing context provide soundness evidence.

### v4

@@ 1 Introduction
However, readability is an aspect of serialization that has been neglected in previous work. Producing readable serialized objects is the problem we address in this paper. This is of utmost importance when serialized objects are consumed by humans. For example, when serialized objects are included in test cases as part of test inputs [10, 11, 12, 13], readability is essential to let developers understand the tests. Clearly, mainstream serialization techniques fall short with respect to readability. For example, 1 shows an example of a test case generated by the automated test generation tool Pankti [12], a tool that creates tests from objects at runtime. In 1, the test inputs are in the XML format, as is the case for the data used to populate the receiver and key variables. XML strings arguably makes the code hard to read, making it challenging to understand the test’s intention.
@@ 15 COSDictionary receiver = new COSDictionary();
To solve the problem of unreadable serialized data, such as JSON/XML test fixtures, we propose the novel concept of plain-code serialization, which consists in serializing an object directly into code statements. By serializing objects to plain code, the serialized representation becomes readable, using a syntax that is both concise and natural to developers. We illustrate this with the semantically equivalent test of 2, which contains the same serialized objects as 1 represented as plain code, instead of XML. In addition, the compiler gives us strong guarantees about the static correctness of the serialized objects. Deserialization then simply means executing the code. In the context of tests, this completely removes dirty data like XML strings, thus making the generated tests more readable and self-contained [17].
@@ 2.1 Concept & Overview
reconstruct(x)=evaluate(plain_code_serialize(x))reconstruct(x)=evaluate(plain\_code\_serialize(x)) (2)
The equivalence relationship ($\equiv$) means that the object after reconstruction is in the same state as the one observed at runtime. While Equation 1 applies to any serialization library, Equation 2 is specific to plain-code serialization, with reconstructreconstruct being the evaluation of the plain-code serialized representation of $x$ obtained from the function $plain\_\allowbreak{}code\_\allowbreak{}serialize$.
@@
1public class Monkey {
2 private int age;
3 public Color eyeColor;
4 public Habitat habitat;
@@
8 public Monkey(int age, Color eyeColor, Habitat habitat) {
9 this(age);
10 this.eyeColor = eyeColor;
11 this.habitat = habitat;
@@
14 public void setEyeColor(Color eyeColor) { this.eyeColor = eyeColor; }
@@
16 public String eatBanana(int dozens) {
17 int quantity = dozens * 12;
18 return Eating + quantity + bananas”;
@@
24public class Habitat {
25 private String coordinates;
26 private double area;
@@
28 public Habitat(String coordinates) { this.coordinates = coordinates; }
@@
30 public void grow(double newArea) {
31 double newValue = this.area + newArea;
32 this.area = newValue;
@@ 2.2 Structure-based Serialization (sb)
The core idea of the structure-based object serialization strategy is to statically analyze a class and build a reconstruction plan. This is done by searching the class for statement which set the value of fields, such as constructor or setter method calls. The plan is instantiated at runtime for a given object.
@@ 2.2.2 Optimizing Readability
As is the case for Monkey, we enumerate all possible reconstruction plans for a target class. However, we are interested in the shortest and most idiomatic plan, to reconstruct an object using the minimum number of statements [18]. To quantify how idiomatic a reconstruction plan is, we first assign each action a pre-defined integer cost value. For example, calling a constructor has a very low cost, while directly setting a field has a higher cost. We then select the reconstruction plan with the lowest cumulative cost. We formulate this optimization as a constraint problem, shown in Equation 4. In the constructed problem, every action $i$ is assigned a boolean selection variable $a_{i}$. If the model value for $a_{i}$ is true, the action is selected. If it is false, the action will not be part of the reconstruction plan. If an action sets a field $f$, it is added to the set $\beta_{f}$, and if it is a constructing action, also to $\alpha$.
The cost of an action $i$ is denoted by $cost_{i}$ and the global optimization objective is modeled with Equation 4a. Equation 4b ensures that at least one constructing action is selected; Equation 4c ensures that there is a single constructing action; and Equation 4d means that all fields are assigned at least once. If we find a solution to the constraint problem, all used actions are deterministically combined to produce a reconstruction plan. The plan will be also applied repeatedly for recursive structures.
$\underset{\displaystyle a\in\{0,1\}^{n}}{\mathrm{min}}\quad\sum_{i=1}^{n}a_{i}\cdot cost_{i}$ (4a)
@@
$a_{1}\lor a_{2}\lor\ldots\lor a_{i}\quad$ $a_{1},\ldots,a_{i}\in\alpha,$ (4b)
$\overline{a_{1}\land a_{2}}$ $a_{1},a_{2}\in\alpha,a_{1}\neq a_{2},$ (4c)
$a_{1}\lor a_{2}\lor\ldots\lor a_{i}\quad$ $f\in fields,a_{1},\ldots,a_{i}\in\beta_{f}$ (4d)
@@ 2.3.1 Capturing Events
An event of interest is the occurrence of an operation which mutates an object. The considered events are constructor calls (ConstructEvent), field assignments (FieldSetEvent), or impure method calls updating a field. Method call events are split into a start (MethodStartEvent) and an end event (MethodEndEvent), allowing other events to happen in-between. We capture all calls to methods with at least one field assignment.
We denote all potential events to reconstruct an object $o$ as
event_sequence(o)event\_sequence(o). Based on this notation, we specialize the general object reconstruction formulation, Equation 1, for trace-based reconstruction as follows:
@@
Given a type we want to serialize, we first identify all types needed for its reconstruction because they are used either as fields, constructor parameters, or method parameters. Then, we instrument the constructors, accessible methods, and field assignments for all these types. For example, in the Habitat class (3), we instrument the constructor on line $28$, as well as the grow method on lines 303330-33. We show the resulting instrumented code for the grow method in 6. When the instrumented grow is invoked at runtime, it emits multiple events, ordered by a monotonically increasing counter representing logical time.
@@ 2.3.3 Object Reconstruction Mixing
We show such a case in 8, where we are constructing a Monkey. Monkey is serialized using structure-based serialization, and its habitat field is serialized using trace-based serialization. As the captured events are mapped to code offline, but the plain-code serialized habitat field has to be inserted into a specific spot in the generated code, the structure-based serialization strategy uses the “use object reference” (from Table 1) marker action. This action emits an id@logicaltimeid@logical-time reference, which is replaced by the trace-based serialized value after the event sequence conversion is finished. Besides this action, Table 1 also contains two more actions, “use static field” and “use standard charset”. These are actions that can not be done structurally, but also do not need the full trace-based serialization machinery. For the “use static field” action type, we look at all static fields of the class we want to serialize and detect if any of these store a reference-equal object. If such a field can be found, we emit a field read instead of tracing the constructor and subsequent mutating method calls. A similar, but more specific, optimization is realized by the “use standard charset” action type. This action type replaces references to Java’s standard charsets with the canonical named constants, and serves as an example of integrating domain logic into the serialization process. To further facilitate extensibility, such custom adapters and action types can be provided as plugins.
@@ 2.4.4 Implementation
ProDJ relies on Spoon [19] for source code analysis, in order to identify the associated types of methods involving serialization within the target Java application. ProDJ uses Z3 [20] to solve the constraint problem of Equation 4 and determine the most idiomatic reconstruction plan for types that are structure-based serializable. The event sequences for trace-based serializable objects are logged in JSON. The instrumentation of methods and their associated types is achieved using ByteBuddy [^3] https://bytebuddy.net/.
@@ 3.2 Subjects
Table 2: Projects used in the evaluation of plain code serialization: the exact version (SHA) we use for our experiments, the number of lines of Java code (LOC), the number of serialization points we consider (#SPs), the number of types related to the serialization points (#Types), the number of serialization points triggered in production by our workloads (#SPs Prod), and the number of types that we target for plain code serialization at runtime (#Types Prod).
We evaluate ProDJ with $4$ open-source Java projects that are popular and active, and span diverse application domains. We present them in Table 2. PDFBox is a library and command-line tool for the manipulation of PDF documents [^5] https://pdfbox.apache.org/. Its features include conversion between text and PDF documents, extraction of images, and merging of PDF documents, among others. GraphHopper is an open-source routing application [^6] https://www.graphhopper.com/, used to compute a path between input locations on a map. BroadLeaf is an e-commerce web application [^7] https://www.broadleafcommerce.com/. Gephi is a graphical tool for working with graph data [^8] https://gephi.org/, including capabilities for visualization and analysis.
For each project in Table 2, column $2$ links to the exact version (SHA) we use for our evaluation, and column $3$ presents the number of lines of Java source code within each project, calculated with tokei. Columns $4$ and $5$ give the number of methods instrumented by ProDJ within the project (#SPs for ‘serialization point’), and consequently, the number of associated types (#Types) that need to be instrumented at this point. For example, ProDJ instruments $2,933$ methods within BroadLeaf, involving $1,883$ types. For our evaluation, we select and instrument methods within each project that are public, non-abstract, non-static, non-deprecated, contain more than one statement, and are declared in a public, non-anonymous, non-local, and non-deprecated class. Columns $6$ and $7$ of Table 2 provide the same metrics, but computed after runtime observations, i.e., the number of serialization points and types that are covered at runtime, which we discuss in the next subsection.
@@ 3.3 Workloads
The key reason we select the projects presented in Table 2 for evaluating plain-code test generation with ProDJ is that we can design representative and repeatable production workloads for each of them, which is essential for our systematic experimental protocols. We invoke PDFBox through its command line interface, and use its “render” command to convert a PDF document containing multiple pages into JPEG image files. We run the GraphHopper server, and query GraphHopper for a route between two coordinates in Berlin [^9] https://bit.ly/3LA2PVu. In BroadLeaf, we add some products to our shopping cart, create a user account, log in, and favorite some products. Next, we browse through our shopping cart and change the quantity of some products, before finally completing our order. We use the graphical interface of Gephi to load the “Les Misérables” dataset [^10] https://github.com/gephi/gephi/wiki/Datasets#social-networks, and compute all the statistics on it that Gephi offers. This includes computing the average degree of nodes, the network diameter, as well as operations like connected components or Pagerank.
Our interactions with each project result in the invocation of serialization points and correspondingly, the instantiation of types that are instrumented. The last two columns of Table 2 give those numbers. For example, of the $2,933$ methods that are instrumented by ProDJ for BroadLeaf, $508$ are triggered by our workload (column #SPs Prod). These methods become the methods under test. Consequently, $444$ unique types involved in serialization are instantiated (column #Types Prod). As we shall see in subsection 4.1, these serialization points will trigger the serialization of thousands of objects under our workload.
@@ 4 Experimental Results
Table 3: Results of plain-code serialization on 44 open-source, real-world Java applications
@@ 4.1 RQ1 (Objects)
ProDJ successfully plain-code serializes thousands of objects ($174,699$) observed at runtime in 44 real-world Java applications. Both the structure-based and trace-based serialization strategies are useful to translate runtime objects in plain Java.
@@
2GHRequest object = new GHRequest();
3// GHPoint uses the structure-based strategy
4GHPoint points = new GHPoint(52.565499, 13.469925);
5GHPoint points0 = new GHPoint(52.508281, 13.326416);
6// The point list uses the structure-based strategy
7List<GHPoint> points1 = new ArrayList<>(List.of(points, points0));
8object.setPoints(points1);
@@ 4.2 RQ2 (Accuracy)
5 AngleCalc receiver = new AngleCalc();
6 double baseOrientation = -1.971212484045175;
7 double orientation = 2.744264992492817;
@@ 5 AngleCalc receiver = new AngleCalc();
9 double actual = receiver.alignOrientation(baseOrientation, orientation);
@@ 15 COSArray receiver = new COSArray();
20 Collection<COSBase> objectsList2 = new ArrayList<>(List.of(objectsList, objectsList0, objectsList1));
@@
In 10, we present one test generated by ProDJ for GraphHopper, clearly following the Arrange-Act-Assert pattern. This test verifies the behavior of the alignOrientation method of the AngleCalc class in GraphHopper. ProDJ generates the Java statements from lines $5$ to $7$ at runtime using the structure-based strategy. Note that ProDJ reuses the parameter names of alignOrientation for the local variables in the test on lines $6$ and $7$. The assertion on line $11$ verifies that the double value output from the invocation of alignOrientation on line $9$ matches the value recorded at runtime. We also present one ProDJ-generated test for PDFBox in 11. The method under test, called getComponents defined in the class PDColor, returns a non-null, float array output, based on the value of a field called colorSpace. Note that ProDJ outlines the recreation of the receiver object of type PDColor into a separate createPDColor method. Within this method, ProDJ uses object reconstruction mixing for plain-code serializing PDColor. Specifically, during the pre-execution phase, ProDJ discovers that the associated type COSArray is eligible for trace-based serialization. At runtime, ProDJ records a constructor call event for COSArray (recreated on line $15$), followed by a call to the method addAll (line $22$). The only parameter of addAll is a list, which ProDJ serializes using the structure-based strategy (line $20$). However, the contents of this list are serialized using the trace-based strategy (lines $16$ - $18$). ProDJ uses the parameter name objectList of addAll to derive the names of the local variables in the helper method. Note that ProDJ observes the calls to the factory method COSFloat.get at runtime and therefore recreates these calls, instead of invoking the COSBase constructor directly. The String arguments for these COSFloat.get method calls are directly inlined, in contrast with the double arguments on lines $6$ and $7$ in the first test of 10. This is because ProDJ tries to inline primitive values, but only if this would not break up the Arrange-Act-Assert pattern. Next, on line $23$ of 11, a PDColorSpace object is reconstructed using the runtime trace as being an instance of the RGB color space. Finally, the two-argument constructor call on line $24$ reuses the reconstructed objects to recreate the runtime PDColor object. This object is used within the generated test as the receiver for the invocation of getComponents (line $8$). Note also the inline creation of the float array on line $10$ within the assertion, which verifies the equality of the expected and actual float array output. We argue that these plain-code serialization and readability optimization strategies employed by ProDJ result in tests that not only reflect production observations, but are also focused, well-structured, an easy-to-read. Each generated test focuses on the invocation of a single target method on the reconstructed receiver object. Consequently, the tests generated by ProDJ do not contain test smells: they are not ‘eager tests’ and do not contain multiple assertions [23]. We note that these tests may miss inconsistencies within other fields of the receiver that are not visible from the assertion.
@@
All test cases generated by ProDJ compile correctly and when we execute them, $1,038$ test cases pass, which is a success rate of 59.959.9%. Specifically, for three of the four case studies, a large majority of the generated tests pass. For PDFBox, 8282% of all tests generated pass, the highest among the four projects. That is a strong property. Given the structure of generated tests, it means that the state of an object after calling a method, i.e., after the ‘Act’ phase of the test, is consistent with the state observed in production (the ‘Assert’ phase). While not being a correctness proof, this is strong consistency evidence.
@@
Our experiment with using plain-code serialization for testing shows that ProDJ faithfully recreates runtime states for all objects in nearly 60% of the generated tests. The assertions in this testing context provide soundness evidence.